A boy tosses a coin upward with a velocity of 14.7 m/s then catches it on his hand. Find the maximum height reached by the coin.


1
Expert's answer
2021-09-09T10:49:31-0400

"h=\frac{v^2}{2g}=\frac{14.7^2}{2\cdot9.8}=11.025\ (m)." Answer